The product of two consecutive positive even intergers is 14 more than their sum. Set up an equation that can be used to find the two numbers and solve it.
Answer by ankor@dixie-net.com(22740)

The product of two consecutive positive even integers is 14 more than their sum.
let x & (x+2) = the two consecutive positive even integers
then
x(x+2) = x + (x+2) + 14
x^2 + 2x = 2x + 16
:
Set up an equation that can be used to find the two numbers and solve it.
Form a quadratic equation on the left
x^2 + 2x - 2x - 16 = 0
x^2 - 16 = 0
Can be factored as the difference of squares
(x-4)(x+4) = 0
The positive solution
x = 4; the next integer is 6
:
:
Check this in the original statement
" The product of two consecutive positive even intergers is 14 more than their sum. "
4 * 6 = 4 + 6 + 14
24 = 10 + 14